# The CLAS Collaboration

### v2

@@ VI.3 Systematic uncertainties
The presence of quasi-elastic events in the dataset facilitates the verification of both the overall cross section normalization and the quality of the electron selection. The former may lack accuracy due to potential miscalibrations of the Faraday cup, fluctuations in the target density, and imprecision knowledge of other parameters involved in the luminosity calculation (see Eq. (10)). The quality of the electron selection in turn may suffer from potential miscalibrations of different detector parts, inaccuracies in the electron tracking and identification, and uncertainties of the cuts and corrections involved in the electron selection.
@@
To verify the cross section normalization and the quality of the electron selection, the quasi-elastic cross section was estimated and compared with the Bosted parameterization of the quasi-elastic cross section off the deuteron Bosted et al. ; Bosted and Christy (2008). This comparison indicates a better than 5% agreement between the experimental and parameterized cross sections and, therefore, a 5% global uncertainty was assigned to the extracted cross sections to account for potential inaccuracies in the normalization and electron selection Skorodumina et al. (2020); Skorodumina .
@@
In this study, the cross sections were extracted in three sets of kinematic variables, as described in Sec. IV.3. The fully integrated cross sections were found to slightly differ among the sets due to different data and efficiency propagation to various kinematic grids. As a final result, integral cross sections averaged (as arithmetic mean) over these three grids are reported and the standard error of the mean is interpreted as a systematic uncertainty Skorodumina et al. (2020); Skorodumina . Since different variable sets correspond to different registered final hadrons (and, therefore, to different combinations of the hadron cuts), the uncertainty due to integration over the three sets of final hadronic variables includes uncertainties due to the shapes of the hadron cuts that were used in the analysis. The average value of this uncertainty among all $W$ and $Q^{2}$ bins is 1.6%.
@@
The cut on the relative efficiency uncertainty performed in this study excluded from consideration all multi-dimensional cells with uncertainties greater than 30% (see Sec. IV.6). To estimate the systematic effect of this cut, the fully integrated cross sections were also calculated for the cut values 25% and 35%. As a final result, the arithmetic mean of the integral cross sections for these three cut values is reported, and the standard error of the mean is interpreted as a systematic uncertainty Skorodumina et al. (2020); Skorodumina . The systematic effect of the relative efficiency uncertainty cut was found to be minor, i.e. the average uncertainty value among all $W$ and $Q^{2}$ bins is 0.8%.
@@
One more part of the systematic inaccuracies comes from the effective correction due to FSI-background admixture. This correction was performed for the experimental events in the $\pi^{-}$ missing topology as described in Sec. III.5.4. The fit shown in Fig. 7 and the corresponding correction factor given by Eq. (5) were found to be slightly dependent on the histogram binning. To account for this uncertainty, the correction factor was calculated for five different histogram bin sizes, and the arithmetic mean of these five individual values was used for the correction (for each bin in $W$). The absolute uncertainty of the resulting correction factor was calculated as the standard error of the mean. The corresponding cross section uncertainty was estimated according to the standard error propagation rules Skorodumina et al. (2020); Skorodumina . The systematic effect of the FSI-background correction was estimated for each bin in $W$ and $Q^{2}$ where the correction was applied. For such bins, the average value of the relative systematic uncertainty is 0.4%.
@@
As a common practice in studies of the double-pion cross sections with CLAS Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018), a 5% global uncertainty was assigned to the cross section due to the inclusive radiative correction procedure (see Sec. V.2).
@@ VI.4 Summary for the cross section uncertainties
Finally, the model-dependent uncertainties $\delta^{\text{cells}}_{\text{model}}(\Delta X)$ and $\delta^{\text{unfold}}_{\text{model}}(\Delta X)$ defined by Eq. (23) and Eq. (25), respectively, were combined with the total statistical uncertainty $\delta_{\text{stat}}^{\text{tot}}(\Delta X)$ defined in Sec. VI.1 as follows,
@@
The extracted cross sections are reported with the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$, which for the single-differential distributions is given by Eq. (26). For the fully integrated cross sections, $\delta_{\text{stat,mod}}^{\text{tot}}$ is obtained from the uncertainty of the single-differential distributions according to the standard error propagation rules. For the majority of integral $(W,Q^{2})$ points, the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ stays on a level of 4%-6%.
For the fully integrated cross sections, in addition to the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$, the total systematic uncertainty is also reported as a separate quantity. If necessary, the relative systematic uncertainty ($\varepsilon_{\text{sys}}^{\text{tot}}$) in each $W$ and $Q^{2}$ bin can be propagated as a global factor to the corresponding single-differential distributions.
For most points of the fully integrated cross sections, uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ is less than the total systematic uncertainty, exceeding the latter only for $W\lesssim 1.4$ GeV. This is because $\delta_{\text{stat,mod}}^{\text{tot}}$ rises near the threshold due to small experimental statistics, large contribution of the empty cells (see Sec. V.1), and pronounced impact of the unfolding correction (see Sec. V.3).
@@
VII Final state interactions
@@ VII.1 FSIs for $\gamma_{v}p(n)\rightarrow p^{\prime}(n^{\prime})\pi^{+}\pi^{-}$
Hadrons produced in exclusive reactions are subject to final state interactions (FSIs). The nature of this phenomenon is complicated due to numerous mechanisms being involved, most of which are driven by the strong interaction Darwish et al. (2003); Tarasov et al. (2011).
@@
The reaction final hadrons are produced in one vertex and after the production, they fly apart in radial directions. The spectator neutron, which was not involved in the production of the final hadrons, is located aside from the production vertex, so that the final hadrons can scatter off the neutron. Therefore, for the double-pion production off protons bound in deuterium, FSIs with the spectator correspond to a combination of proton-neutron Shirokov and Yudin (1980); Rohrlich and Eisenstein (1949) and pion-neutron scattering Cutkosky et al. (1979); Gasparyan et al. (2003); Vrana et al. (2000), and thus represent a superposition of a broad spectrum of mechanisms inherent for these two scattering types.
Due to the relatively low energy of the final hadrons in this experiment, the majority of FSIs in the investigated reaction are thought to happen elastically, which implies that the quantum numbers of the participating hadrons do not change and no new particles are produced in such interactions Shirokov and Yudin (1980).
@@ VII.2 Distortions due to FSIs
The two FSI types introduced above have some kinematical distinctions from each other and their impact on the extracted cross sections also differ.
An important feature of FSIs among the final hadrons is that they preserve the net four-momentum of all three final hadrons (as long as no new particles are produced in such interactions). For this reason, this FSI type does not affect missing mass distributions Skorodumina . However, interactions among the final hadrons still alter the momenta of the individual particles, and thus introduce distortions into the measured cross sections (no matter whether off a free or bound nucleon). These cross section distortions can hardly be avoided on the level of the experimental data analysis due to the insensitivity of the missing mass distributions to this FSI type. This issue needs to be accounted for at the level of theoretical/phenomenological cross section interpretation.
Meanwhile, interactions with the spectator nucleon have one distinctive difference from FSIs among the final hadrons. Specifically, as the spectator nucleon is extrinsic to the original exclusive reaction, any interaction with it changes the total four-momentum of the reaction final state. As a consequence, events in which final hadrons interacted with the spectator, introduce distortions into missing mass distributions Skorodumina et al. (2021). These distortions reveal agglomerations of FSI-affected events, which allows for their separation from the quasi-free event sample, so that the quasi-free cross sections can be extracted (see more details in Sec. III.5).
In this analysis, distortions due to FSIs with the spectator neutron are clearly visible in experimental distributions of the missing quantities $P_{X}$ and $M^{2}_{X[\pi^{-}]}$, where the former is defined by Eq.(2) and the latter is $[P_{X[\pi^{-}]}^{\mu}]^{2}$ with the $P_{X[\pi^{-}]}^{\mu}$ defined by Eq.(3). These distortions were found to differ in different regions of the reaction phase space and also to be topology dependent. Illustrations and further details are given in the following subsections.
Figure 11: $M^{2}_{X[\pi^{-}]}$ distributions for the fully exclusive (left) and for the $\pi^{-}$ missing (right) topologies in five 100-MeV-wide bins in $W$. The mismatch between the experimental (black) and the simulated (blue) histograms reveals agglomerations of events in which the final hadron interacted with the spectator neutron. The distributions are normalized in a way that the peak maxima are equal to one and then zoomed in on the range [0, 0.25] on the $y$-axis, to better visualize the mismatch. The presented statistics correspond to the unzoomed experimental distributions.
@@ VII.3 Comparison between the two topologies
Figure 11 presents the $M^{2}_{X[\pi^{-}]}$ distributions for the fully exclusive (left) and the $\pi^{-}$ missing (right) topologies in five 100-MeV-wide bins in $W$. Experimental distributions are shown in black, while blue histograms correspond to simulated distributions of pure quasi-free events. The mismatch between them therefore reveals agglomerations of events in which the final hadrons interacted with the spectator neutron. The Monte Carlo simulation was performed on the basis of the TWOPEG-D event generator Skorodumina et al. (2017a), which successfully reproduces the Fermi smearing of the missing quantities, but does not include FSI effects. The distributions are normalized in a way that the peak maxima are equal to one and then zoomed in on the range [0, 0.25] on the $y$-axis, to better visualize the mismatch.
Final hadrons attributed to various topologies have different kinematics and therefore different probabilities to interact with the spectator neutron. For this reason, events affected by FSIs with the spectator are distributed differently in the two reaction topologies as illustrated by Fig. 11. More details can be found in Ref. Skorodumina .
@@ VII.4 FSIs in the fully exclusive topology
In general, the proportion of events affected by FSIs with the spectator was found to vary greatly (from negligible to considerable) in different ranges of final hadron momenta/angles.
Remarkably, some regions of the reaction phase space were found to be completely dominated by quasi-free events. For example, Fig. 12 shows a good match between the experimental (black) and the simulated (blue) distributions of the quantities $P_{X}$ (left) and $M^{2}_{X[\pi^{-}]}$ (right) for $\pi^{-}$ momentum magnitudes from 0.5 GeV to 1.4 GeV, which indicates the dominance of quasi-free events in this momentum range.
Conversely, other regions of the reaction phase space were revealed to be mostly populated by events with FSI-disturbed kinematics. Figure 13 illustrates this effect for proton polar angles ranging from ${}^{\circ}$ to ${}^{\circ}$, where a large mismatch between the experimental (black) and the simulated (blue) distributions is observed, which reveals a considerable fraction of events affected by FSI with the spectator in this kinematic region. More illustrations can be found in Ref. Skorodumina .
@@
Figure 13: Experimental (black) and simulated (blue) distributions of the quantities $P_{X}$ (left) and $M^{2}_{X[\pi^{-}]}$ (right) for the proton polar angle ranging from ${}^{\circ}$ to ${}^{\circ}$. The distributions are plotted for the fully exclusive topology and normalized in a way that the maxima of the main peaks are equal to one.
@@ VII.5 FSIs in topologies with a missing hadron
In topologies with an unregistered hadron $i$, the quantity $M^{2}_{X[i]}$ is typically used for the channel identification (it is defined analogously to $M^{2}_{X[\pi^{-}]}$ introduced above). For reactions occurring off bound nucleons, interactions between the final hadrons and the spectator nucleon have a different impact on $M^{2}_{X[i]}$ depending on which final hadron experienced the FSI.
@@
    All final hadrons in an event avoided interactions with the neutron. Then this event is a true quasi-free event and the four-momentum of the unregistered hadron can be successfully reconstructed as missing.
@@
    The unregistered hadron avoided FSIs, while one of the registered hadrons interacted with the neutron, changing its four-momentum and hence losing its kinematic affiliation to the initial reaction. This does not allow for proper reconstruction of the missing hadron four-momentum, causing the event to contribute to the FSI-background in the $M^{2}_{X[i]}$ distributions Skorodumina et al. (2021).
@@
    The unregistered hadron interacted with the neutron and the registered hadrons did not. In this case, the missing four-momentum of the unregistered hadron corresponds to its four-momentum before the FSI. Such an event then kinematically mimics a quasi-free event.
@@
This disposition reveals that for reactions off bound nucleons, topologies with a missing hadron suffer from the presence of falsely defined quasi-free events, which are events of the third type. Such events are kinematically indistinguishable from true quasi-free events and for this reason this effect can hardly be corrected for.
VIII Extracted quasi-free cross sections
[…]

### v4

@@ VI.3 Systematic uncertainties
The presence of quasielastic events in the dataset facilitates the verification of both the overall cross section normalization and the quality of electron selection. The former may lack accuracy due to potential miscalibrations of the Faraday cup, fluctuations in the target density, and imprecise knowledge of other parameters involved in the luminosity calculation [see Eq. (10)]. The electron selection quality in turn may suffer from potential miscalibrations of different detector parts, inaccuracies in electron tracking and identification, and uncertainties of the cuts and corrections involved in the electron selection.
@@
To verify the cross section normalization and the quality of the electron selection, the quasielastic cross section was estimated and compared with the Bosted parameterization of the quasielastic cross section off the deuteron Bosted et al. ; Bosted and Christy (2008). This comparison indicated a better than 5% agreement between the experimental and the parameterized cross sections, and therefore a 5% global uncertainty was assigned to the extracted double-pion cross sections to account for potential inaccuracies in the normalization and electron selection Skorodumina et al. (2020); Skorodumina .
@@
In this study, the cross sections were extracted in three sets of kinematic variables, as described in Sec. IV.3. The fully integrated cross sections $\sigma_{\text{v}}^{\text{int}}$ were found to slightly differ among the sets due to different data and efficiency propagation to various kinematic grids. For this reason, the arithmetic mean of the cross sections $\sigma_{\text{v}}^{\text{int}}$ obtained in the three variable sets is reported as a final result, and the standard error of the mean is interpreted as a systematic uncertainty Skorodumina et al. (2020); Skorodumina . Since different variable sets correspond to different registered final hadrons (and, therefore, to different combinations of the hadron cuts), this uncertainty includes potential inaccuracies in the hadron identification. The average value of this uncertainty among all $W$ and $Q^{2}$ bins is 1.6%.
@@
The cut on the relative efficiency uncertainty performed in this study excluded from consideration all multi-dimensional cells with uncertainties greater than 30% (see Sec. IV.6). To estimate the systematic effect of this cut, the fully integrated cross sections were also calculated for the cut values 25% and 35%. As a final result, the arithmetic mean of the cross sections $\sigma_{\text{v}}^{\text{int}}$ calculated for these three cut values is reported, and the standard error of the mean is interpreted as a systematic uncertainty Skorodumina et al. (2020); Skorodumina . The systematic effect of the relative efficiency uncertainty cut was estimated for each bin in $W$ and $Q^{2}$, and the average uncertainty value among all bins was found to be 0.8%.
@@
One more part of the systematic inaccuracies comes from the effective correction due to the FSI-background admixture. This correction was performed for experimental events in the $\pi^{-}$ missing topology as described in Sec. III.5.4. The fit shown in Fig. 7 and the corresponding correction factor given by Eq. (5) were found to be slightly dependent on the histogram binning. To account for this uncertainty, the correction factor was calculated for five different histogram bin sizes, and the arithmetic mean of these five individual values was used for the correction. The absolute uncertainty of the resulting correction factor was calculated as the standard error of the mean. The corresponding cross section uncertainty was estimated according to the standard error propagation rules Skorodumina et al. (2020); Skorodumina . The systematic uncertainty due to the FSI-background correction was estimated for each bin in $W$ and $Q^{2}$ where the correction was applied, and its average value was found to be 0.4%.
@@
As a common practice in studies of double-pion cross sections with CLAS Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018), a 5% global uncertainty was assigned to the cross sections due to the inclusive radiative correction procedure (see Sec. V.2).
@@ VI.4 Summary for the cross section uncertainties
Finally, the model-dependent uncertainties $\delta^{\text{cells}}_{\text{model}}(\Delta X)$ and $\delta^{\text{unfold}}_{\text{model}}(\Delta X)$ defined by Eqs. (23) and (25), respectively, were combined with the total statistical uncertainty $\delta_{\text{stat}}^{\text{tot}}(\Delta X)$ defined in Sec. VI.1 as follows:
@@
The extracted cross sections are reported with the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$, which for the single-differential distributions is given by Eq. (26). For the fully integrated cross sections, $\delta_{\text{stat,mod}}^{\text{tot}}$ is obtained from the uncertainty of the single-differential distributions according to the standard error propagation rules. For the majority of integral $(W,Q^{2})$ points, the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ stays on a level of 4-6%.
For the fully integrated cross sections, in addition to the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$, the total systematic uncertainty is also reported as a separate quantity. If necessary, the relative systematic uncertainty $\varepsilon_{\text{sys}}^{\text{tot}}$ can be propagated as a global factor to the corresponding single-differential distributions in each $W$ and $Q^{2}$ bin.
For most points of the fully integrated cross sections, the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ is less than the total systematic uncertainty, exceeding the latter only for $W\lesssim 1.4$ GeV. This is because $\delta_{\text{stat,mod}}^{\text{tot}}$ rises near the reaction threshold due to small experimental statistics, large contributions from the empty cells (see Sec. V.1), and pronounced impact of the unfolding correction (see Sec. V.3).
@@
VII Final-state interactions
@@ VII.1 FSIs for $\gamma_{v}p(n)\rightarrow p^{\prime}(n^{\prime})\pi^{+}\pi^{-}$
Hadrons produced in exclusive reactions are subject to final-state interactions (FSIs). The nature of this phenomenon is complicated due to numerous mechanisms being involved, most of which are driven by the strong interaction Darwish et al. (2003); Tarasov et al. (2011).
@@
The reaction final hadrons are produced in one vertex, and after the production, they fly apart in radial directions. The spectator neutron, which was not involved in the reaction, is located aside from the production vertex, so that the final hadrons can scatter off the neutron. Therefore, for double-pion production off protons bound in deuterium, FSIs with the spectator correspond to a combination of proton-neutron Shirokov and Yudin ; Rohrlich and Eisenstein (1949) and pion-neutron scattering Cutkosky et al. (1979); Gasparyan et al. (2003); Vrana et al. (2000) and thus represent a superposition of a broad spectrum of mechanisms inherent for these two scattering types.
Due to the relatively low energy of the final hadrons in this experiment, the majority of FSIs in the investigated reaction are thought to happen elastically, which implies that the quantum numbers of the participating hadrons do not change, and no new particles are produced in such interactions Shirokov and Yudin .
@@ VII.2 Distortions due to FSIs
The two FSI types introduced above have some kinematic distinctions from each other and their impacts on the extracted cross sections also differ.
An important feature of FSIs among the final hadrons is that they preserve the net four-momentum of all three final hadrons (as long as no new particles are produced in such interactions). For this reason, this FSI type does not affect missing mass distributions Skorodumina . However, interactions among the final hadrons still alter the momenta of the individual particles and thus introduce distortions into the measured cross sections (no matter whether off a free or bound nucleon). These cross section distortions can hardly be avoided at the level of experimental data analysis due to insensitivity of missing mass distributions to this FSI type. This issue needs to be accounted for at the level of theoretical and/or phenomenological cross section interpretation.
Meanwhile, interactions with the spectator nucleon have one distinctive difference from FSIs among the final hadrons. Specifically, as the spectator nucleon is extrinsic to the original exclusive reaction, any interaction with it changes the total four-momentum of the reaction final state. As a consequence, events in which the final hadrons interacted with the spectator introduce distortions into missing mass distributions Skorodumina et al. (2021). The distortions reveal agglomerations of FSI-affected events, which allows for their separation from the quasifree event sample, so that the quasifree cross sections can be extracted (see more details in Sec. III.5).
In this analysis, distortions due to FSIs with the spectator neutron are clearly visible in experimental distributions of the missing quantities $P_{X}$ and $M^{2}_{X[\pi^{-}]}$, where the former is defined by Eq.(2) and the latter is $[P_{X[\pi^{-}]}^{\mu}]^{2}$ with the $P_{X[\pi^{-}]}^{\mu}$ defined by Eq.(3). These distortions were found to differ in different regions of the reaction phase space and also to be topology dependent. Illustrations and further details are given in the following sections.
Figure 11: $M^{2}_{X[\pi^{-}]}$ distributions for the fully exclusive (left) and the $\pi^{-}$ missing (right) topologies in five 100-MeV-wide bins in $W$. The mismatch between the experimental (black) and the simulated (blue) histograms reveals agglomerations of events in which the final hadrons interacted with the spectator neutron. The distributions are normalized in a way that the peak maxima are equal to one and then zoomed in on the range [0, 0.25] on the $y$ axis, to better visualize the mismatch. The presented statistics correspond to the unzoomed experimental distributions.
@@ VII.3 Comparison between the two topologies
Figure 11 presents the $M^{2}_{X[\pi^{-}]}$ distributions for the fully exclusive (left) and the $\pi^{-}$ missing (right) topologies in five 100-MeV-wide bins in $W$. Experimental distributions are shown in black, while the blue histograms correspond to simulated distributions of pure quasifree events. The mismatch between them therefore reveals agglomerations of events in which the final hadrons interacted with the spectator neutron. The Monte Carlo simulation was performed on the basis of the TWOPEG-D event generator Skorodumina et al. (2017a), which successfully reproduces the Fermi smearing of the missing quantities but does not include FSI effects. The distributions are normalized in a way that the peak maxima are equal to one and then zoomed in on the range [0, 0.25] on the $y$ axis, to better visualize the mismatch.
Final-state hadrons attributed to various topologies have different kinematics and therefore different probabilities to interact with the spectator neutron. For this reason, events affected by FSIs with the spectator are distributed differently in the two reaction topologies as illustrated by Fig. 11. More details can be found in Ref. Skorodumina .
@@ VII.4 FSIs in the fully exclusive topology
In general, the proportion of events affected by FSIs with the spectator was found to vary greatly (from negligible to considerable) in different ranges of the final hadron momenta and/or angles.
Remarkably, some regions of the reaction phase space were found to be completely dominated by quasifree events. For example, Fig. 12 shows a good match between the experimental (black) and the simulated (blue) distributions of the quantities $P_{X}$ (left) and $M^{2}_{X[\pi^{-}]}$ (right) for the $\pi^{-}$ momentum magnitude from 0.5 to 1.4 GeV, which indicates the dominance of quasifree events in this momentum range.
Figure 13: Experimental (black) and simulated (blue) distributions of the quantities $P_{X}$ (left) and $M^{2}_{X[\pi^{-}]}$ (right) for the proton polar angle ranging from 40 to 60. The distributions are plotted for the fully exclusive topology and normalized in a way that the maxima of the main peaks are equal to one.
Conversely, other regions of the reaction phase space were revealed to be mostly populated with events in FSI-disturbed kinematics. Figure 13 illustrates this effect for the proton polar angle range from 40 to 60, where a large mismatch between the experimental (black) and the simulated (blue) distributions is observed, which reveals a considerable fraction of events affected by FSIs with the spectator in this kinematic region. More illustrations can be found in Ref. Skorodumina .
@@ VII.5 FSIs in topologies with a missing hadron
In topologies with an unregistered hadron $i$, the quantity $M^{2}_{X[i]}$ is typically used for the channel identification (it is defined analogously to $M^{2}_{X[\pi^{-}]}$ introduced above). For reactions occurring off bound nucleons, interactions between the final hadrons and the spectator nucleon have different impacts on $M^{2}_{X[i]}$ depending on which final hadron experienced the FSI.
@@
    All final hadrons in an event avoided interactions with the neutron. Then this event is a true quasifree event and the four-momentum of the unregistered hadron can be successfully reconstructed as missing.
@@
    The unregistered hadron avoided FSIs, while one of the registered hadrons interacted with the neutron, changing its four-momentum and hence losing kinematic affiliation to the original exclusive reaction. This does not allow for proper reconstruction of the missing hadron four-momentum, causing the event to contribute to the FSI background in the $M^{2}_{X[i]}$ distributions Skorodumina et al. (2021).
@@
    The unregistered hadron interacted with the neutron and the registered hadrons did not. In this case, the missing four-momentum of the unregistered hadron corresponds to its four-momentum before the FSI. Such an event then kinematically mimics a quasifree event.
@@
This disposition reveals that for reactions off bound nucleons, topologies with a missing hadron suffer from the presence of falsely defined quasifree events, which are events of the third type. Such events are kinematically indistinguishable from true quasifree events and for this reason this effect can hardly be corrected for.
VIII Extracted quasifree cross sections
Figure 15: Single-differential cross sections for $W=1.5375$ GeV and $Q^{2}=0.625$ GeV2. The error bars correspond to the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ defined in Sec. VI.4. The cross section estimation shown by the solid curves is based on the free proton event generator TWOPEG Skorodumina et al. (2017b) (see text for more details).